# definition

### v3

@@ 1 Introduction
Heterogeneity poses a fundamental challenge for many real-world decision-making problems, where each problem consists of a large number of interacting components. However, despite its practical significance, heterogeneity remains largely understudied in the literature. In this paper, we study heterogeneous settings of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs) [23]. A WCMDP consists of $N$ arms (or subproblems), where each arm itself is a Markov decision process (MDP). In a heterogeneous setting, the MDPs could be distinct. At each time step, the decision-maker selects an action for each arm, which affects the arm’s transition probabilities and reward, and then the arms make state transitions independently. However, these actions are subject to a set of global budget constraints, where each constraint limits one type of total cost across all arms at each time step. The objective is to find a policy that maximizes the long-run average reward over an infinite time horizon. We focus on the planning setting, where all the model parameters (reward function, cost functions, budget and transition probabilities) are known.
@@
Although heterogeneity is crucial in the modeling of these applications, most existing work on average-reward WCMDPs (or their special cases) establish asymptotic optimality only for the homogeneous setting where all arms share the same set of model parameters [42, 39, 14, 15, 25, 26, 27, 47, 24, 22]. The only known exceptions are [39] and [46]. They consider the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N$ scales up, with each type having distinct model parameters. Their budget constraints are also of special forms. While heterogeneous WCMDPs have been studied under the finite-horizon total-reward and discounted-reward criteria, as we review at the end of this section and also in Appendix A, these results do not extend to the average-reward setting we consider.
@@
We first focus on the average-reward criterion. As mentioned earlier, most existing work considers the homogeneous setting. Early work on WCMDPs primarily focuses on a special case known as the restless bandit (RB) problem, where each arm’s MDP has a binary action space (active and passive actions) and there is only one budget constraint that limits the total number of active actions across all arms at each time step. The seminal work by Whittle [44] introduced the RB problem and the celebrated Whittle index policy, which was later shown to achieve an $o(1)$ optimality gap as $N\to\infty$ under a set of conditions [42]. Subsequent work on RBs has focused on designing policies that achieve asymptotic optimality under more relaxed conditions [39, 25, 26, 47], as well as improving the optimality gap to $O(1/\sqrt{N})$ [25, 26] or $O(\exp(-cN))$ [14, 15, 27]. Among these papers, only [39] addresses heterogeneous RBs, but in the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N\to\infty$.
@@ 2 Problem setup
Our objective is to maximize the long-run time-average reward subject to the cost constraints. To be more precise, we follow the treatment for maximizing average reward in [35]. For any policy $\pi$ and an initial state $\bm{S}_{0}$ of the $N$-armed system, consider the limsup average reward $R^{+}(\pi,\bm{S}_{0})$ and the liminf average $R^{-}(\pi,\bm{S}_{0})$, defined as $R^{+}(\pi,\bm{S}_{0})=\lim\sup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$ and $R^{-}(\pi,\bm{S}_{0})=\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$. If $R^{+}(\pi,\bm{S}_{0})=R^{-}(\pi,\bm{S}_{0})$, then the average reward of policy $\pi$ under initial condition $\bm{S}_{0}$ exists and is defined as
@@
Our LP (5) serves a similar role to the LP used in previous work on restless bandits and WCMDPs with homogeneous arms [see, e.g., 42, 14, 25]. Both our LP and the LP in previous work relax the hard budget constraints to time-average constraints. However, in the homogeneous arm setting, the LP has only one set of state-action frequencies $y(s,a)$, and the LP is independent of $N$. As a result, both the optimal value of the LP and the complexity of solving it are independent of $N$. Some existing work [39] considers heterogeneous arms, but only in the limited sense of having a constant number of arm types. This setting still closely resembles the homogeneous setting, and the LP remains independent of $N$.
In contrast, our work addresses fully heterogeneous arms. Consequently, we must define separate state-action frequencies $y_{i}(s,a)$ for each arm $i\in[N]$, making our LP explicitly depend on $N$. Therefore, the optimal value $R_{N}^{\textup{rel}}$ depends on $N$, and the complexity of solving the LP grows with $N$. Nevertheless, because the number of variables and constraints scale linearly with $N$, our LP can still be solved in polynomial time.
@@ Appendix A Detailed review on related work
In this section, we provide a more detailed (but still non-exhaustive) review of the literature. We mainly focus on theoretical work with formal performance guarantees, leaving out the extensive body of work with empirical results. We first survey papers with the same reward criterion as ours, i.e., infinite-horizon average-reward criterion. In this setting, we begin with papers on homogeneous restless bandits (RBs), which is a special case considered by most existing papers. Then we give a more detailed review of the papers on average-reward WCMDPs. Afterward, we review the papers on WCMDPs with two other reward criteria, i.e., the finite-horizon total-reward criterion and the infinite-horizon discounted-reward criterion. Finally, we briefly mention other problems that are related to WCMDPs.
Infinite-horizon average-reward homogeneous RBs.
@@
The first asymptotic optimality result for average-reward homogeneous RBs was established by Weber and Weiss [42]: it was shown that the Whittle index policy [44] achieves an $o(1)$ optimality gap as the number of arms $N$ goes to infinity. There are three key assumptions in [42]: indexability, the global attractor property, and the aperiodic-unichain condition. These assumptions are gradually relaxed in the subsequent papers. In particular, [39] proposed a class of priority policies based on an LP relaxation. This class of policies, later referred to as the LP-Priority policies, generalizes the Whittle index policy. Each LP-Priority policy achieves an $o(1)$ optimality gap without requiring indexability. Later, Hong et al. [25, 26] introduced new policies that further removed the global attractor property and improved the optimality gap to $O\bigl{(}1/\sqrt{N}\bigr{)}$. More recently, Yan [47] proposed the align-and-steer policy, which further weakened the aperiodic-unichain condition and achieved an $o(1)$ optimality gap.
@@
Parallel to relaxing the assumptions for asymptotic optimality, another of line of work has focused on improving the optimality gap beyond $O(1/\sqrt{N})$ under slightly stronger assumptions [14, 15, 27]. Specifically, Gast et al. [15] showed that the Whittle index policy has an $O(\exp(-cN))$ optimality gap for some constant $c>0$. In addition to indexability and the aperiodic-unichain condition, [15] also requires a stronger version of global attractor property named Uniform Global Attractor Property (UGAP), and a condition called non-singularity. Subsequently, Gast et al. [14] showed that LP-Priority policies achieve $O(\exp(-cN))$ optimality gaps assuming the aperiodic-unichain condition, UGAP, and a non-degenerate condition that is equivalent to non-singularity. More recently, Hong et al. [27] proposed a two-set policy that also achieves an $O(\exp(-cN))$ optimality gap while replacing UGAP of [14] with a much weaker condition named local stability.
@@
The papers on average-reward WCMDPs remain relatively scarce, and to our knowledge, fully heterogeneous WCMDPs have yet to be addressed. Nevertheless, some papers consider special cases of WCMDPs that generalize restless bandits by allowing multiple actions, more general constraints, or typed heterogeneity. In particular, Hodge and Glazebrook [24] generalized Whittle index policy to homogeneous WCMDPs with a single-constraint and multiple actions, where each action represents a different activation level and has a different cost. Verloop [39] extended the LP-Priority policies to typed heterogeneous WCMDPs with a single constraint and multiple actions, but requires each action to have the same cost function. Then Xiong et al. [46] proposed another index policy for typed heterogeneous WCMDPs with a single constraint, and allows each action to have a different cost. In the three papers above [24, 39, 46], $o(1)$ optimality gaps have been proved under a similar set of assumptions as in most restless bandit papers, i.e., aperiodic-unichain (or irreducibility) condition, global attractor property, and a generalized indexability condition if the policy is Whittle index. More recently, Goldsztajn and Avrachenkov [22] considered homogeneous WCMDPs and proposed a class of policies with $o(1)$ optimality gaps under a weaker-than-standard aperiodic-unichain condition.
@@
Next, we review the asymptotic optimality results for finite-horizon total-reward RBs and WCMDPs. The finite-horizon setting is better understood than the average-reward setting, partly because the analysis in the finite horizon is not hindered by the technical conditions arising from average-reward MDPs, such as the unichain condition and the global attractor property. On the other hand, computing asymptotically optimal policies for the finite-horizon setting is more complicated, requiring a careful optimization of the transient sample paths.
@@
Hu and Frazier [28] proposed the first asymptotic optimal policy for finite-horizon homogeneous RBs, which achieves an $o(1)$ optimality gap without any assumptions. [^1] Here, we measure the optimality gap in terms of the reward per arm, to be consistent with our convention. However, in the papers on the finite-horizon total-reward setting, it is also common to measure the optimality gap in terms of the total reward of all arms, which differs from ours by a factor of $N$. We also stick to the same convention when reviewing the papers on the infinite-horizon discounted-reward setting. Since then, researchers have established asymptotic optimality in more general settings [49, 12, 17, 8, 10]. Among these papers, the most general setting was addressed by Brown and Zhang [10], which considered fully heterogeneous WCMDPs; they obtained $O(1/\sqrt{N})$ optimality gaps for a naive fluid policy and a reoptimization-based fluid policy, among a few other results to be reviewed in the next paragraphs. Notably, there is also a further generalization of fully heterogeneous WCMDPs, which involves an exogenous state that affects all arms’ transitions, rewards, and constraints; Brown and Zhang [9] proposed this setting, where they achieved an $O(1/\sqrt{N})$ optimality gap using a dynamic fluid policy.
@@
Another line of papers improved the optimality gap beyond the order $O(1/\sqrt{N})$ by making an additional assumption called non-degeneracy. Specifically, Zhang and Frazier [51] established an $O(1/N)$ optimality gap in non-degenerate homogeneous RBs. Gast et al. [14] then proposed a different policy for the same setting that improved the optimality gap to $O(\exp(-cN))$. Later, Gast et al. [16] and Brown and Zhang [10] established $O(1/N)$ optimality gaps for homogeneous and typed heterogeneous WCMDPs, respectively, assuming non-degeneracy. More recently, Zhang [50] proposed a policy for fully heterogeneous WCMDPs; the optimality gap bound of the policy interpolates between $O(1/\sqrt{N})$ and $O(1/N)$ as the degree of non-degeneracy varies, unifying the performance bounds in the degenerate and non-degenerate worlds.
@@
Despite the generality of the settings and the fast diminishing rate of the optimality gaps as $N\to\infty$, most of the optimality gaps in the finite-horizon setting depend super-linearly on the time horizon, so they do not carry over to the infinite-horizon average-reward setting. There are two exceptions, [10, 16], which achieve optimality gaps that depend linearly on the time horizon under some special conditions: [10] requires all entries of the transition kernels to be bounded away from zero; [16] assumes an ergodicity property, which requires two arms in any different states synchronize in a fixed number of steps under any sequence of actions with a positive probability. However, without these conditions, the optimality gaps in [10, 16] depend quadratically on the time horizon. Apart from having distinct optimality gap bounds, all existing algorithms in the finite-horizon setting need to (sometimes repeatedly) solve LPs whose number of variables scale with the time horizon, so they cannot be directly adapted to the infinite-horizon average-reward setting.
@@
Asymptotic optimality has also been established for RBs and WCMDPs under the infinite-horizon discounted-reward criterion. In particular, Brown and Smith [8] established an $O(N^{\log_{2}(\sqrt{\gamma})})$ optimality gap for fully heterogeneous WCMDPs when $\gamma\in(1/2,1)$. Subsequently, Zhang and Frazier [52], Ghosh et al. [17] obtained $O(1/\sqrt{N})$ optimality gaps for homogeneous and typed heterogeneous RBs, and Brown and Zhang [10] established the same order of optimality gap for fully heterogeneous WCMDPs. Similar to the finite-horizon setting, most of these optimality gaps depend super-linearly on the effective time horizons $1/(1-\gamma)$ unless special conditions hold [10], so they do not carry over to the infinite-horizon average-reward setting. The policies here also require solving LPs whose complexities scale with the effective time horizon.

### v4

@@ 1 Introduction
Heterogeneity poses a fundamental challenge for many real-world decision-making problems, where each problem consists of a large number of interacting components. However, despite its practical significance, heterogeneity remains largely understudied in the literature. In this paper, we study heterogeneous settings of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs) [23]. A WCMDP consists of $N$ arms (or subproblems), where each arm itself is a Markov decision process (MDP). In a heterogeneous setting, the MDPs could be distinct. At each time step, the decision-maker selects an action for each arm, which affects the arm’s transition probabilities and reward, and then the arms make state transitions independently. However, these actions are subject to a set of global budget constraints, where each constraint limits one type of total cost across all arms at each time step. The objective is to find a policy that maximizes the long-run average reward over an infinite time horizon. We focus on the planning setting, where all the model parameters (reward function, cost functions, budget, and transition probabilities) are known.
@@
Although heterogeneity is crucial in the modeling of these applications, most existing work on average-reward WCMDPs establishes asymptotic optimality only for the homogeneous setting where all arms share the same set of model parameters [42, 39, 15, 14, 25, 26, 27, 47, 24, 22]. Only a few exceptions [39, 46, 25] address heterogeneity, but in highly specialized settings. Among these, a common approach to handle heterogeneity is to consider the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N$ scales up, with each type having distinct model parameters. While heterogeneous WCMDPs have been studied under the finite-horizon total-reward and discounted-reward criteria, these results do not extend to the average-reward setting we consider. We review related work in more detail at the end of this section and also in Appendix A.
@@
We first focus on the average-reward criterion. As mentioned earlier, most existing work considers the homogeneous setting. Early work on WCMDPs primarily focuses on a special case known as the restless bandit (RB) problem, where each arm’s MDP has a binary action space (active and passive actions) and there is only one budget constraint that limits the total number of active actions across all arms at each time step. The seminal work by Whittle [44] introduced the RB problem and the celebrated Whittle index policy, which was later shown to achieve an $o(1)$ optimality gap as $N\to\infty$ under a set of conditions [42]. Subsequent work on RBs has focused on designing policies that achieve asymptotic optimality under more relaxed conditions [39, 25, 26, 47], as well as improving the optimality gap to $O(1/\sqrt{N})$ [25, 26] or $O(\exp(-cN))$ [15, 14, 27]. Among these papers, [39, 25] address heterogeneous RBs. However, [39] focuses on the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N\to\infty$. The paper [25] includes an extension to the fully heterogeneous setting. However, for their result to yield asymptotic optimality, there need to be further assumptions on the orders of the so-called synchronization times in the paper. The policies in both papers cannot be straightforwardly extended to general WCMDPs, which have multiple actions, multiple cost constraints, and state-dependent cost functions.
@@ 2 Problem setup
Our objective is to maximize the long-run time-average reward subject to the cost constraints. To be more precise, we follow the treatment for maximizing average reward in [35]. For any policy $\pi$ and an initial state $\bm{S}_{0}$ of the $N$-armed system, consider the limsup average reward $R^{+}(\pi,\bm{S}_{0})$ and the liminf average $R^{-}(\pi,\bm{S}_{0})$, defined as $R^{+}(\pi,\bm{S}_{0})=\lim\sup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$ and $R^{-}(\pi,\bm{S}_{0})=\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$. If $R^{+}(\pi,\bm{S}_{0})=R^{-}(\pi,\bm{S}_{0})$, then the average reward of policy $\pi$ under the initial condition $\bm{S}_{0}$ exists and is defined as
@@
Our LP (5) serves a similar role to the LP used in previous work on restless bandits and WCMDPs (see, e.g., [42, 39, 15, 24, 46, 22, 25]), but with different forms and dependencies on $N$. Both our LP and the LP in previous work relax the hard budget constraints to time-average constraints. However, in the homogeneous arm setting [42, 15, 24, 22], the LP has only one set of state-action frequencies $y(s,a)$, and the LP is independent of $N$. As a result, both the optimal value of the LP and the complexity of solving it are independent of $N$. Prior work for the typed-heterogeneous setting [46, 39, 25] divides the arms into a constant number $K$ of types and defines a set of state-action frequency $y_{k}(s,a)$ for each type $k$. The optimal value of the resulting LP and the complexity of solving it depend on the number of types $K$. The work [25] includes a generalization to the fully heterogeneous setting, resulting in an LP similar to ours, although for the special case of restless bandits.
In our LP, we define a separate set of state-action frequencies $y_{i}(s,a)$ for each arm $i\in[N]$, making the LP explicitly depend on $N$. Therefore, the optimal value $R_{N}^{\textup{rel}}$ depends on $N$, and the complexity of solving the LP grows with $N$. Nevertheless, because the number of variables and constraints scales linearly with $N$, our LP can still be solved in polynomial time.
@@ Appendix A Detailed review on related work
In this section, we provide a more detailed, though still non-exhaustive, review of the literature. We mainly focus on theoretical work with formal performance guarantees, leaving out the extensive body of work with empirical results. We begin by surveying papers with the same reward criterion as ours, i.e., infinite-horizon average-reward criterion. In this setting, we first review existing work on restless bandits (RBs), which is an extensively studied special case of WCMDPs. We then give a more detailed review of existing results on WCMDPs. Next, we turn to work that considers other reward criteria: the finite-horizon total-reward criterion and the infinite-horizon discounted-reward criterion. Finally, we briefly mention other problems that are related to WCMDPs.
Infinite-horizon average-reward RBs.
@@
For the homogeneous arm setting, the first asymptotic optimality result for average-reward homogeneous RBs is established by Weber and Weiss [42]: they show that the Whittle index policy [44] achieves an $o(1)$ optimality gap as the number of arms $N$ goes to infinity. There are three key assumptions in [42]: indexability, the global attractor property, and the aperiodic-unichain condition. These assumptions have been gradually relaxed in the subsequent papers. In particular, [39] proposes a class of priority policies based on an LP relaxation. This class of policies, later referred to as the LP-Priority policies, generalizes the Whittle index policy. Each LP-Priority policy achieves an $o(1)$ optimality gap without requiring indexability. The work [25] is the first one that breaks the global attractor property assumption. The authors propose a policy named Follow-the-Virtual-Advice (FTVA), which achieves an $O(1/\sqrt{N})$ optimality gap under an assumption named the Synchronization Assumption; there exist problem instances that satisfy the Synchronization Assumption but do not satisfy the global attractor property. Later work [26] further relaxes the conditions and only requires the aperiodic-unichain condition to achieve an $O(1/\sqrt{N})$ optimality gap. More recently, Yan [47] proposes the align-and-steer policy, which further weakens the aperiodic-unichain condition and achieves an $o(1)$ optimality gap.
@@
Parallel to relaxing the assumptions for asymptotic optimality, another line of work has focused on improving the optimality gap beyond $O(1/\sqrt{N})$ under slightly stronger assumptions [15, 14, 27]. Specifically, Gast et al. [14] show that the Whittle index policy has an $O(\exp(-cN))$ optimality gap for some constant $c>0$. In addition to indexability and the aperiodic-unichain condition, [14] also requires a stronger version of the global attractor property named Uniform Global Attractor Property (UGAP), and a condition called non-singularity. Subsequently, Gast et al. [15] show that LP-Priority policies achieve $O(\exp(-cN))$ optimality gaps assuming the aperiodic-unichain condition, UGAP, and a non-degenerate condition that is equivalent to non-singularity. More recently, Hong et al. [27] propose a two-set policy that also achieves an $O(\exp(-cN))$ optimality gap while replacing UGAP of [15] with a much weaker condition named local stability.
Among the aforementioned work, [39] and [25] have addressed the heterogeneous arm setting. The setting studied in [39] is the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N\to\infty$. The paper [25] includes an extension to the fully heterogeneous setting. In particular, the proposed FTVA policy generalizes to the fully heterogeneous setting and leads to an optimality gap of $O(\overline{\tau}^{\text{sync}}_{\max}/\sqrt{N})$, where $\overline{\tau}^{\text{sync}}_{\max}$ is the maximum of a quantity called the synchronization time across all arms. Therefore, for this result to yield asymptotic optimality, there needs to be a further assumption that $\overline{\tau}^{\text{sync}}_{\max}=o(\sqrt{N})$. We re-emphasize that the FTVA policy does not generalize to WCMDPs. The main reason is that FTVA heavily relies on the fact that an RB only constrains the number of pulls, while a WCMDP has budget constraints on cost functions each depending on both the state and action of an arm.
@@
Work on average-reward WCMDPs remains relatively scarce, and to our knowledge, fully heterogeneous WCMDPs have yet to be addressed. Compared to RBs, a WCMDP allows multiple actions for each arm and multiple cost constraints, where each cost function is a function of both the state and the action of an arm. The line of research [24, 39, 46, 22] has generalized the action space and cost model of RBs to WCMDPs, and some of them allow for typed heterogeneity. In particular, Hodge and Glazebrook [24] generalize the Whittle index policy to homogeneous WCMDPs with a single constraint and multiple actions, where each action represents a different activation level and has a different cost. Verloop [39] extends the LP-Priority policies to typed heterogeneous WCMDPs with a single constraint and multiple actions, but requires each action to have the same cost. Then Xiong et al. [46] propose another index policy for typed heterogeneous WCMDPs with a single constraint, and allow each action to have a different cost. In the three papers above [24, 39, 46], $o(1)$ optimality gaps have been proved under a similar set of assumptions as in most restless bandit papers, i.e., aperiodic-unichain (or irreducibility) condition, global attractor property, and a generalized indexability condition if the policy is Whittle index. Finally, Goldsztajn and Avrachenkov [22] consider homogeneous WCMDPs with multiple actions and multiple cost constraints with general cost functions, and propose a class of policies with $o(1)$ optimality gaps under a weaker-than-standard aperiodic-unichain condition.
@@
Next, we review the asymptotic optimality results for finite-horizon total-reward RBs and WCMDPs. The finite-horizon setting is better understood than the average-reward setting, partly because the analysis in the finite horizon is not hindered by the technical conditions arising in average-reward MDPs, such as the unichain condition and the global attractor property. On the other hand, computing asymptotically optimal policies for the finite-horizon setting is more complicated, requiring a careful optimization of the transient sample paths.
@@
Hu and Frazier [28] propose the first asymptotically optimal policy for finite-horizon homogeneous RBs, which achieves an $o(1)$ optimality gap without any assumptions. [^1] Here, we measure the optimality gap in terms of the reward per arm, to be consistent with our convention. However, in the papers on the finite-horizon total-reward setting, it is also common to measure the optimality gap in terms of the total reward of all arms, which differs from ours by a factor of $N$. We also adopt the same convention when reviewing the papers on the infinite-horizon discounted-reward setting. Since then, researchers have established asymptotic optimality in more general settings [49, 12, 17, 8, 10]. Among these papers, the most general setting is addressed by Brown and Zhang [10], where the authors consider fully heterogeneous WCMDPs; they obtain $O(1/\sqrt{N})$ optimality gaps for a naive fluid policy and a reoptimization-based fluid policy, among a few other results to be reviewed in the next paragraphs. Notably, there is also a further generalization of fully heterogeneous WCMDPs, which involves an exogenous state that affects all arms’ transitions, rewards, and constraints; Brown and Zhang [9] propose this setting, where they achieve an $O(1/\sqrt{N})$ optimality gap using a dynamic fluid policy.
@@
Another line of work has improved the optimality gap beyond the order $O(1/\sqrt{N})$ by making an additional assumption called non-degeneracy. Specifically, Zhang and Frazier [51] establish an $O(1/N)$ optimality gap in non-degenerate homogeneous RBs. Gast et al. [15] then propose a different policy for the same setting that improves the optimality gap to $O(\exp(-cN))$. Later, Gast et al. [16] and Brown and Zhang [10] establish $O(1/N)$ optimality gaps for homogeneous and typed heterogeneous WCMDPs, respectively, assuming non-degeneracy. More recently, Zhang [50] proposes a policy for fully heterogeneous WCMDPs; the optimality gap bound of the policy interpolates between $O(1/\sqrt{N})$ and $O(1/N)$ as the degree of non-degeneracy varies, unifying the performance bounds in the degenerate and non-degenerate cases.
@@
Despite the generality of the settings and the fast diminishing rate of the optimality gaps as $N\to\infty$, most of the optimality gaps in the finite-horizon setting depend super-linearly on the time horizon, so they do not carry over to the infinite-horizon average-reward setting. There are two exceptions, [10, 16], which achieve optimality gaps that depend linearly on the time horizon under some special conditions: [10] requires all entries of the transition kernels to be bounded away from zero; [16] assumes an ergodicity property, which requires two arms in any different states to synchronize in a fixed number of steps under any sequence of actions with a positive probability. However, without these conditions, the optimality gaps in [10, 16] depend quadratically on the time horizon. Apart from having distinct optimality gap bounds, all existing algorithms in the finite-horizon setting need to (sometimes repeatedly) solve LPs whose number of variables scales with the time horizon, so they cannot be directly adapted to the infinite-horizon average-reward setting.
@@
Asymptotic optimality has also been established for RBs and WCMDPs under the infinite-horizon discounted-reward criterion. In particular, Brown and Smith [8] establish an $O(N^{\log_{2}(\sqrt{\gamma})})$ optimality gap for fully heterogeneous WCMDPs when $\gamma\in(1/2,1)$. Subsequently, Zhang and Frazier [52], Ghosh et al. [17] obtain $O(1/\sqrt{N})$ optimality gaps for homogeneous and typed heterogeneous RBs, and Brown and Zhang [10] establish the same order of optimality gap for fully heterogeneous WCMDPs. Similar to the finite-horizon setting, most of these optimality gaps depend super-linearly on the effective time horizon $1/(1-\gamma)$ unless special conditions hold [10], so they do not carry over to the infinite-horizon average-reward setting. The policies here also require solving LPs whose complexities scale with the effective time horizon.